write an equation in slope intercept form of the line having the givin slope and containing given point
m:3/2, (4,6)

Answer by mananth(16949)

m= -3/ 2

Plug value of the slope and point (4,6) in
The equation of any line in slope intercept form is given by
Y = m x + b
Plug value of the slope and point (4,6) in the above equation to get b

6.00 = -6 + b
b= 6 - (-6 )
b= 12
So the equation will be
Y = -3/2 x + 12
m.ananth@hotmail.ca